OpenStudy (anonymous):

The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle?

OpenStudy (zephyr141):

perimeter=2L+2W ---------------- 68=2(W-6)+2W 68=2W-12+2W 56=4W 14=W ---------------- 68=2L+2(14) 40=2L 20=L i'm just throwing numbers around. i hope someone comes by and sees if i did this right or wrong because i feel wrong about it. haha.